Calculating the future value of property. Ben Collins plans to buy a house for $220,000. If...

  1. Calculating the future value of property. Ben Collins plans to buy a house for $220,000. If that real estate is expected to increase in value by 3 percent each year, what will its approximate value be seven year from now?

  2. Using the Rule of 72. Using the rule of 72, approximate the following amounts:
    1. If the value of land in an area is increasing 6 percent a year, how long will it take for property values to double?
    2. If you earn 10 percent on your investments, how long will it take for your money to double?
    3. At an annual interest rate of 5 percent, how long will it take for your savings to double?

  3. Determining the inflation rate. In 2006, selected automobiles had an average cost of $16,000. The average cost of those same automobiles is now $28,000. What was the rate of increase for these automobiles between the two time periods?

  4. Computing the Time Value of Money. Using time value f money tables, calculate the following:
    1. The future value of $450 six years from now at 7 percent.
    2. The future value of $900 saved each year for 10 years at 8 percent.
    3. The amount a person would have to deposit today (present value) at a 6 percent interest rate to have $1,000 five years from now.
    4. The amount a person would have to deposit today to be able to take out $600 a year for 10 years from an account earning 8 percent.

  5. Calculating the Present Value of a Series. Pete Morton is planning to go to graduate school in a program of study that will take three years. Pete wants to have $15,000 available each year for various school and living expenses. If he earns 4 percent on his money, how much must be deposited at the start of his studies to be able to withdraw $15,000 a year for three years?

  6. Determining a Loan Payment Amount. If you borrow $8,000 with a 5 percent interest rate, to be repaid in five equal yearly payments, what would be the amount of each payment? (Note: Use the present value of an annuity table.)
Answer #1

1.
=220000*1.03^7=270572.250393471

2.
=72/6=12

3.
=72/10=7.2

4.
=72/5=14.4

5.
=(28000/16000)^(1/(2009-2006))-1=20.5071132087615%

6.
=450*1.07^6=675.32865833205

7.
=900/8%*(1.08^10-1)=13037.9062193189

8.
=1000/1.06^5=747.258172866057

9.
=600/8%*(1-1/1.08^10)=4026.04883936487

10.
=15000/4%*(1-1/1.04^3)=41626.3654984069

11.
=8000*5%/(1-1/1.05^5)=1847.79838502614